# Sampling Theory for Function Approximation with Numerical Redundancy

### v1

@@ Abstract
The study of numerical rounding errors is often greatly simplified in the analytical treatment of mathematical problems, or even entirely separated from it. In sampling theory, for instance, it is standard to assume the availability of an orthonormal basis for computations, ensuring that numerical errors are negligible. In reality, however, this assumption is often unmet. In this paper, we discard it and demonstrate the advantages of integrating numerical insights more deeply into sampling theory. To clearly pinpoint when the numerical phenomena play a significant role, we introduce the concept of numerical redundancy. A set of functions is numerically redundant if it spans a lower-dimensional space when analysed numerically rather than analytically. This property makes it generally impossible to compute the best approximation of a function in its span using finite precision. In contrast, $\ell^{2}$-regularized approximations are computable and, therefore, form the foundation of many practical methods. Regularization generally reduces accuracy compared to the best approximation, but our analysis shows that there is a benefit: it also significantly reduces the amount of data needed for accurate approximation. In this paper, we develop the analytical tools needed to fully understand this effect. Furthermore, we present a constructive method for optimally selecting data points for $L^{2}$-approximations, explicitly accounting for the effects of regularization. The results are illustrated for two common scenarios that lead to numerical redundancy: (1) approximations on irregular domains and (2) approximations that incorporate specific features of the function to be approximated. In doing so, we obtain new results on random sampling for Fourier extension frames. Finally, we establish that regularization is implicit in numerical orthogonalization of a numerically redundant set, indicating that its analysis can not be bypassed in a much broader range of methods.
@@ 1 Introduction
We are interested in the problem of computing an approximation to a function $f$ in a Hilbert space $H$. Due to the focus on computability, this problem will be referred to as numerical approximation. An important aspect is that although the function $f$ generally resides in an infinite-dimensional space, a numerical method can process only a finite number of measurements of $f$. This naturally leads to the introduction of a sequence of finite-dimensional spaces $\{V_{n}\}_{n\in\mathbb{N}}$ of increasing dimension $n$, forming approximations to $H$. It is hereby assumed, based on prior knowledge, that $f$ is well approximated in the spaces $V_{n}$ in the sense that fast convergence is attainable. Associated with a numerical approximation is thus an operator $\mathcal{F}_{n,m}:\mathbb{C}^{m}\to V_{n}$ mapping $m$ measurements of a function $f$ to a function $f_{n,m}$ in the approximation space $V_{n}$. Note that the number of measurements $m$ can differ from the dimension $n$. In particular, it is often required that $m$ is sufficiently large compared to $n$, or equivalently $n$ sufficiently small, in order to obtain robust and accurate approximation methods [3, 18].
Another important aspect of numerical approximation is how to represent an approximation in $V_{n}$. The computed approximations $f_{n,m}$ are continuous objects, which cannot be stored directly on a computer. To this end, one needs a spanning set $\Phi_{n}=\{\phi_{i,n}\}_{i=1}^{n^{\prime}}$ satisfying $\operatorname{span}(\Phi_{n})=V_{n}$. This allows the approximation $f_{n,m}$ to be characterized by a set of expansion coefficients $\mathbf{x}$
@@
Sampling theory aims to determine the amount of information required from a function $f$ to achieve a (near-)optimal approximation in $V_{n}$, while ensuring robustness to noise. Standard results for linear approximation methods are typically obtained by analysing $\mathcal{F}_{n,m}$ and, hence, are independent of the spanning set $\Phi_{n}$ that is used. The underlying assumption is that the numerical computations are performed using an orthonormal or Riesz basis as a spanning set for $V_{n}$, such that $\mathcal{L}_{n,m}$ and $\mathcal{F}_{n,m}$ have similar properties. However, this assumption is often not satisfied in practice; many different types of spanning sets show up in a wide variety of applications. Riesz bases for $V_{n}$ are often unknown and computing them can be numerically unstable. Furthermore, the choice of the spanning set may be influenced by the potential to reduce the computational cost of the method; structured spanning sets can permit the use of fast transforms such as the Fast Fourier Transform.
In particular, many applications make use of a numerically redundant spanning set. Such a set might be linearly independent from an analytic point of view, but linearly dependent from a numerical point of view. In this case, regularization is needed to obtain a bounded map $\mathcal{L}_{n,m}$ from measurements to expansion coefficients. In [5, 6], this phenomenon is analysed for regularization by truncation of the singular value decomposition (TSVD). It is shown that the approximation space after regularization is effectively smaller than $V_{n}$ and, therefore, regularization generally lowers the accuracy. On the other hand, it is also demonstrated that the number of required measurements reduces due to regularization, depending on the spanning set that is used. Hence, in these cases, the choice of the spanning set does influence the sampling theoretical analysis. The primary focus of this paper is identifying when and how the chosen spanning set impacts the required amount of data for accurate function approximation.
We describe the approximation problem in more detail and give known results about sampling for least squares approximations to frame the main contribution and overview of the paper.
@@ 1.1 Approximation problem
A typical example is the approximation of continuous functions in $H=L^{2}([-1,1])$ using pointwise evaluations. In this case, $G$ is the space of continuous functions $C([-1,1])$ with associated norm $\|g\|_{G}=\sup_{x\in[-1,1]}\lvert g(x)\rvert$ satisfying (3) for $C_{G}=2$. Furthermore, if
@@
and $\mathbf{n}\in\mathbb{C}^{m}$ is measurement noise. The question arises if the approximation is accurate, i.e., if it lies close to the best approximation in $V_{n}$ and is robust to measurement noise. We will first discuss the well-studied case of least squares approximation.
@@ 1.2 Least squares approximation
In this paper, we limit ourselves to error bounds that depend on the optimal error measured in the $G$-norm. Deriving bounds in terms of the $H$-norm is typically more challenging. For instance, probabilistic error bounds of this kind can be obtained when considering random pointwise samples for $H=L^{2}$ approximations, as demonstrated in [19, 23].
@@
It follows from Theorem 1 that the accuracy and robustness of least squares approximations depend solely on $A_{n,m}$, $C_{G}$ and $\|\mathcal{M}_{m}\|_{G,2}$, which in turn do not depend on the chosen spanning set $\Phi_{n}$. However, recall from (6) that one cannot directly compute $\mathcal{F}_{n,m}^{\text{LS}}$, yet computes expansion coefficients via $\mathcal{L}_{n,m}^{\text{LS}}$. Hence, the numerical stability of the algorithm also depends on $\mathcal{L}_{n,m}^{\text{LS}}$, which can behave wildly for certain choices of $\Phi_{n}$. More specifically, it holds for any $\mathbf{d}\in\mathbb{C}^{m}$ that
@@
where $\sigma_{\min}(\cdot)$ denotes the smallest nonzero singular value and $K\coloneq\max_{v\in V_{n},\|v\|_{H}=1}\|v\|_{m}$. As explained in section 2, $\sigma_{\min}(\mathcal{T}_{n})$ is close to zero for numerically redundant spanning sets and, hence, regularization is generally needed to obtain a bounded operator $\mathcal{L}_{n,m}$ mapping measurements to expansion coefficients. The goal of this paper is to deepen our understanding of sampling for numerically redundant spanning sets $\Phi_{n}$, taking into account the effects of regularization.
@@ 1.3 Main contribution
As mentioned in the introduction, the influence of $\ell^{2}$-regularization in the case of a numerically redundant spanning set was first highlighted in [5, 6]. This led to an analysis of the truncated singular value decomposition (TSVD) as a means for regularizing least squares approximations. The accuracy of a TSVD approximation is described in [6, Theorem 1.3] and depends on a regularization parameter, as well as on the existence of accurate approximations with small expansion coefficients. Furthermore, the accuracy depends on two constants, $\kappa^{\epsilon}_{M,N}$ and $\lambda^{\epsilon}_{M,N}$, which characterize the influence of discretization. These constants are influenced by regularization, yet in a non-transparent way.
The primary goal of this paper is to deepen our understanding of sampling for $\ell^{2}$-regularized least squares approximations. Our main contribution is the identification of a new norm inequality, analogous to that of unregularized least squares approximation (9), which fully characterizes the accuracy and conditioning of these methods. The following results are part of Theorem 7 and 8.
@@
A couple of important conclusions are in order. First and foremost, (13) is a relaxation of (9); one can always find $A_{n,m}^{\epsilon}\geq A_{n,m}$. In order to see why, observe that both norm inequalities are equivalent for $\epsilon=0$, i.e., in the case of working with infinite precision. In other words: it is easier to satisfy (13) than (9) and, hence, regularization can only reduce the required amount of data. After formally introducing the concept of numerical redundancy, it will become clear that this relaxation is significant only for numerically redundant spanning sets. Second, the accuracy is now governed by (14) as opposed to (10). As discussed in section 2.2, this behaviour is expected when approximating with finite precision. As a result, the size of the expansion coefficients $\|\mathbf{x}\|_{2}$ is an important object of study, which was also one of the main conclusions of [5, 6]. Third, whereas (9) solely depends on the sampling operator $\mathcal{M}_{m}$ and the approximation space $V_{n}$, (13) also depends on the spanning set $\Phi_{n}$ and the regularization parameter $\epsilon$. These additional dependencies make its analysis significantly more complicated.
These results offer a new perspective on the influence of finite precision. While finite precision requires regularization, which generally reduces accuracy compared to the best approximation, there is a benefit: regularization also decreases the amount of required information. Predicting the exact interplay between these two effects is, however, challenging. The main difficulty lies in the dependency of (13) and (14) on the specific spanning set $\Phi_{n}$. Whereas earlier work has focused on the influence of regularization on the convergence behaviour [5, 6], we develop the tools to understand its influence on the associated sampling problem.
@@ 1.4 Paper Overview
In section 2, we formalize the notion of numerical redundancy using ideas from low rank matrix theory and harmonic analysis. Our primary goal is to clearly identify when the effects discussed in this paper are significant. Using this framework we prove that the achievable error of a numerical approximation deviates from the analytical best approximation error, particularly if the spanning set is numerically redundant. Furthermore, through a backward stability analysis we show that simple least squares fitting generally fails to produce accurate results in the presence of numerical redundancy, while the accuracy of $\ell^{2}$-regularized approximations is close to optimal from a numerical point of view. The results in section 2 also hint at the idea that the effects of finite precision cannot be overcome through straightforward numerical orthogonalization. This is investigated further in section 6, where it is demonstrated that regularization is implicit in numerical orthogonalization of a numerically redundant set. This implies that the results of this paper are relevant for a much broader class of methods.
Due to the necessity of $\ell^{2}$-regularization when computing with numerically redundant spanning sets, it is worthwhile to take regularization into account during discretization. Section 3 formalizes how regularization relaxes the discretization condition, resulting in a decrease of the amount of data required for accurate function approximation. Our analysis is not limited to TSVD regularization, or even $\ell^{2}$-regularization; instead, we present a unified framework for regularized least squares problems with a general penalty term as the underlying methodology seamlessly extends. In section 4, we show an additional advantage of the formulation (13): it enables constructive methods for selecting the sampling functionals $l_{j,m}$. Specifically, we identify an efficient random sampling distribution for $L^{2}$-approximations that takes into account the influence of regularization. Section 5 applies the results to two settings of numerically redundant approximations: approximation that incorporates prior knowledge about the asymptotic behaviour of the function to be approximated and approximation on irregular domains. In doing so, we obtain new results on sampling for Fourier extension frames. Specifically, we show that only a log-linear amount of uniformly random samples is sufficient for convergence down to machine precision in a Fourier extension frame. In contrast, an analysis that ignores the effects of regularization would suggest a need for quadratic oversampling.
2 Approximating with finite precision
In this section, we discuss the influence of working with a finite relative precision, yet we do so before introducing a discretization for the problem at hand. An advantage of this approach is that the results do not depend on the properties of a sampling operator. However, this also means that whereas we gain insight in numerical approximation, the results in this section are not yet numerically computable.
2.1 Numerical redundancy
@@
i.e., $\mathbf{x}$ lies in the kernel of the synthesis operator $\mathcal{T}_{n}$ (5). When computing numerically, the notion of linear dependency changes drastically since computations can only be performed with a finite working precision $\operatorname{\epsilon_{\text{mach}}}$. On a machine with finite precision, a vector $\mathbf{x}$ is stored as $\bar{\mathbf{x}}$ where
@@
$\|\mathcal{T}_{n}\bar{\mathbf{x}}-\mathcal{T}_{n}\mathbf{x}\|_{2}=\|\mathcal{T}_{n}\Delta\mathbf{x}\|_{2}\leq\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}\|\mathbf{x}\|_{2}.$ (16)
@@
To take into account the influence of finite precision, it is customary in low rank matrix theory to introduce a numerical rank. Whereas the rank of $\mathcal{T}_{n}$ equals $n$, its $\epsilon$-rank can be defined analogously to [28] by
@@
where the infimum is taken over all linear operators $\mathcal{X}_{n}$ mapping from $\mathbb{C}^{n^{\prime}}$ to $H$. The $\epsilon$-rank takes into account that, in the presence of rounding errors, the computed output $\mathcal{T}_{n}\bar{\mathbf{x}}$ behaves similarly as the exact output $\mathcal{X}_{n}\mathbf{x}$ of some linear operator $\mathcal{X}_{n}$ that lies close to $\mathcal{T}_{n}$. Indeed, similarly to (16) one has
$\|\mathcal{X}_{n}\mathbf{x}-\mathcal{T}_{n}\mathbf{x}\|_{2}\leq\epsilon\|\mathcal{T}_{n}\|_{2,H}\|\mathbf{x}\|_{2}$
@@
A spanning set $\Phi_{n}$ for the $n$-dimensional space $V_{n}$ is called “numerically redundant” if the $\epsilon$-rank of its synthesis operator $\mathcal{T}_{n}$ is strictly less than $n$, where $\epsilon=\operatorname{\epsilon_{\text{mach}}}$ is the working precision associated with the numerical computations.
An elegant characterization of numerically redundant spanning sets follows from the field of harmonic analysis. To this end we introduce frames, which extend the concept of a Riesz basis to allow for the inclusion of linearly dependent sets. As defined by [17, Definition 1.1.1], the set $\Phi_{n}$ is a frame for $V_{n}$ if there exist frame bounds $0<A_{n}\leq B_{n}$ such that
@@
If we refer to the frame bounds $A_{n}$ and $B_{n}$ of $\Phi_{n}$ we will always assume that they are optimal, meaning that they are the largest and smallest constant that satisfy (18), respectively. $A_{n}$ and $B_{n}$ always exist, since every finite set is a frame for its span [17, Corollary 1.1.3]. Furthermore, they satisfy
@@
From the extremal properties of singular values [13, Lecture 23], it follows that
@@
where the minimum is taken over all linear operators $\mathcal{X}_{n}$ mapping from $\mathbb{C}^{n^{\prime}}$ to $H$. Combining this result with Definition 1, one can conclude the $\Phi_{n}$ is numerically redundant if and only if the smallest nonzero singular value of $\mathcal{T}_{n}$ is less than or equal to $\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}$. The final result follows (19). ∎
@@
Based on the properties of $\Psi$, one can determine whether $\Phi_{n}$ is numerically redundant for sufficiently large $n$. In [5, section 4] this is analysed for a linearly independent infinite-dimensional frame $\Psi$. A set $\Psi$ is an infinite-dimensional frame for $H$ if there exist constants $A,B>0$ such that
@@
Furthermore, $\Psi$ is linearly independent if every finite subsequence is linearly independent [17, §3-4]. Note that in contrast to the finite-dimensional case, it does not hold that every infinite set is a frame for its closed span. Again, if we refer to the frame bounds $A$ and $B$ of $\Psi$ we assume that they are optimal. If $\Psi$ is a frame for $H$, analogously to (20) one has that for any $f\in H$ there exist coefficients $\mathbf{x}\in\ell^{2}(\mathbb{N})$ satisfying
@@
Assuming that $\Psi$ is a linearly independent frame, numerical redundancy can now be identified from analysing the upper and lower frame bound of $\Phi_{n}$ as $n$ increases. Following [5, Lemma 5], the upper frame bound $B_{n}$ behaves nicely; it is shown that $\{B_{n}\}_{n\in\mathbb{N}}$ is monotonically increasing and $B_{n}\to B$, as $n\to\infty$. However, the lower frame bound $A_{n}$ can behave wildly; it is shown that $\{A_{n}\}_{n\in\mathbb{N}}$ is monotonically decreasing and $\inf_{n}A_{n}=0$ if $\Psi$ is overcomplete, which means that there exist unit-norm coefficients $\mathbf{c}\in\ell^{2}(\mathbb{N})$ for which

### v2

@@ Abstract
We explore the interaction between numerical rounding errors and discretization errors involved in linear function approximation schemes. To precisely identify when numerical effects become significant, we introduce the concept of numerical redundancy: a set of functions is numerically redundant if, due to finite-precision arithmetic, it spans a lower-dimensional space numerically than it does analytically. This phenomenon arises in a variety of settings, though it is not always explicitly identified as such. When such a set is used as a basis for numerical approximations, recovering the expansion coefficients of the best approximation is generally impossible. More specifically, we show that approximations produced by practical algorithms—affected by rounding errors—are implicitly subject to $\ell^{2}$-regularization. The computed solutions, therefore, exhibit reduced accuracy compared to the best approximation. On the other hand, there is also a benefit: regularization reduces the amount of data needed for accurate and stable discrete approximation. The primary aim of this paper is to develop the tools needed to fully understand this effect. Additionally, we present a constructive method for optimally selecting random data points for $L^{2}$-approximations, explicitly accounting for the influence of rounding errors. We illustrate our results in two typical scenarios that lead to numerical redundancy: (1) approximation on irregular domains and (2) approximation that incorporate specific features of the function to be approximated. In doing so, we obtain new results on random sampling for Fourier extension frames.
Keywords. function approximation, weighted least squares, regularization, orthogonalization, frames theory
MSC codes. 42C15, 41A65, 65F22, 65T40, 65F25
@@ 1 Introduction
We consider the problem of computing an approximation to a function $f$ in a Hilbert space $H$. Due to the focus on computability, this problem will be referred to as numerical approximation. An important aspect is that although the function $f$ generally resides in an infinite-dimensional space, a numerical method can process only a finite number of measurements of $f$. This naturally leads to the introduction of a sequence of finite-dimensional spaces $\{V_{n}\}_{n\in\mathbb{N}}$ of increasing dimension $n$, forming approximations to $H$. It is hereby assumed, based on prior knowledge, that $f$ is well approximated in the spaces $V_{n}$ in the sense that fast convergence is attainable. Associated with a numerical approximation is thus an operator $\mathcal{F}_{n,m}:\mathbb{C}^{m}\to V_{n}$ mapping $m$ measurements of a function $f$ to a function $f_{n,m}$ in the approximation space $V_{n}$. Note that the number of measurements $m$ can differ from the dimension $n$. In particular, it is often required that $m$ is sufficiently large compared to $n$, or equivalently $n$ sufficiently small, in order to obtain robust and accurate approximation methods [4, 20].
Another important aspect of numerical approximation is how to represent an approximation in $V_{n}$. To this end, one chooses a spanning set $\Phi_{n}=\{\phi_{i,n}\}_{i=1}^{n^{\prime}}$ satisfying $\operatorname{span}(\Phi_{n})=V_{n}$. This allows the approximation $f_{n,m}$ to be characterized by a set of expansion coefficients $\mathbf{x}$
@@
In sampling theory, the central goal is to determine how much information about $f$ is needed to construct a (near-)optimal approximation in $V_{n}$, while ensuring robustness to noise. Classical results for linear approximation are typically obtained by analysing $\mathcal{F}_{n,m}$, and are thus independent of the choice of the spanning set $\Phi_{n}$ for $V_{n}$. The underlying assumption is that the numerical computations are performed using an orthonormal or Riesz basis, such that $\mathcal{L}_{n,m}$ and $\mathcal{F}_{n,m}$ have similar properties.
However, this assumption is often violated. In particular, computational methods frequently use numerically redundant spanning sets $\Phi_{n}$—that is, due to finite-precision arithmetic, the functions in $\Phi_{n}$ span a lower-dimensional space numerically than they do in exact arithmetic. This leads to discretized systems that are numerically rank-deficient, and hence subject to significant rounding errors. In order to understand what happens numerically, the analysis must account for the interaction between rounding and discretization errors. Unlike in classical settings, it is no longer sufficient to only study $\mathcal{F}_{n,m}$; the analysis heavily depends on $\mathcal{L}_{n,m}$ making our results inherently dependent on the choice of representation $\Phi_{n}$, not merely on the approximation space $V_{n}$.
Numerical redundancy arises in a number of settings. For example, in function approximation on irregular domains, orthonormal or Riesz bases are not readily available. A standard strategy in such settings is to restrict a known basis—originally defined on a surrounding, regular domain—to the target approximation domain, yielding a numerically redundant spanning set [40]. Note that the target domain may sometimes be unknown, as is often the case for high-dimensional approximation methods arising in uncertainty quantification [7]. Redundancy also occurs when bases are combined, enriched or weighted to better capture specific features of the function being approximated [31, 45, 33]. Other settings of numerical redundancy include radial basis function approximations [46], Trefftz methods [37], and deep neural networks [3]. The analysis of redundant sets for numerical approximation was initiated in [6, 8].
In the presence of numerical redundancy, the coefficients corresponding to the best approximation are generally not computable. Instead, practical algorithms affected by rounding errors compute approximations that are effectively $\ell^{2}$-regularized. While some numerical methods attempt to circumvent this by orthogonalizing the basis, we demonstrate that standard numerical orthogonalization techniques are not able to mitigate the effects of finite-precision arithmetic. Therefore, our results apply broadly across many numerical approximation schemes that rely on redundant representations.
We describe the approximation problem in more detail and give known results about sampling for least squares approximations before stating the main contribution and overview of the paper.
@@ 1.1 Approximation problem
A typical example is the approximation of continuous functions in $H=L^{2}([-1,1])$ using pointwise evaluations. In this case, $G$ can be the space of continuous functions $C([-1,1])$ with associated norm $\|g\|_{G}=\sup_{x\in[-1,1]}\lvert g(x)\rvert$ satisfying (3) for $C_{G}=2$. Furthermore, if
@@
and $\mathbf{n}\in\mathbb{C}^{m}$ is measurement noise. The question arises whether the approximation is accurate, i.e., whether it lies close to the best approximation in $V_{n}$ and is robust to measurement noise. We will first discuss the well-studied case of least squares approximation.
@@ 1.2 Least squares approximation
In this paper, we limit ourselves to error bounds that depend on the optimal error measured in the $G$-norm. Deriving bounds in terms of the $H$-norm typically requires additional effort. For instance, probabilistic error bounds of this kind can be obtained when considering random pointwise samples for $H=L^{2}$ approximations, as demonstrated in [21, 25].
Remark 3.
@@
It follows from Theorem 1 that the accuracy and robustness of least squares approximations depend solely on $A_{n,m}$, $C_{G}$ and $\|\mathcal{M}_{m}\|_{G,2}$, which in turn do not depend on the chosen spanning set $\Phi_{n}$. However, recall from (6) that one does not directly compute $\mathcal{F}_{n,m}^{\text{LS}}$ yet obtains the expansion coefficients via the mapping $\mathcal{L}_{n,m}^{\text{LS}}$. Although Theorem 1 guarantees accuracy and stability, the norm of $\mathcal{L}_{n,m}^{\text{LS}}$ can be arbitrarily large for certain choices of spanning sets, leading to significant rounding errors. More precisely, for any $\mathbf{d}\in\mathbb{C}^{m}$ we have
@@
where $\sigma_{\min}$ denotes the smallest nonzero singular value and $K\coloneq\max_{v\in V_{n},\|v\|_{H}=1}\|v\|_{m}$. We discuss in section 2 that $\sigma_{\min}(\mathcal{T}_{n})$ is close to zero for numerically redundant spanning sets. Consequently, as elaborated on in section 3, practical algorithms implementing $\mathcal{L}_{n,m}^{\text{LS}}$ are subject to $\ell^{2}$-regularization. The aim of this paper is to deepen our understanding of sampling strategies for numerically redundant spanning sets $\Phi_{n}$, taking into account the associated numerical effects.
@@ 1.3 Main contribution
The role of $\ell^{2}$-regularization in the case of a numerically redundant spanning set was first highlighted in [6, 8]. This led to an analysis of the truncated singular value decomposition (TSVD) as a means for regularizing least squares approximations. The accuracy of a TSVD approximation is described in [8, Theorem 1.3] and depends on a regularization parameter, as well as on the existence of accurate approximations with small expansion coefficients. Furthermore, the accuracy depends on two constants, $\kappa^{\epsilon}_{M,N}$ and $\lambda^{\epsilon}_{M,N}$, which characterize the influence of discretization. These constants are influenced by regularization, yet in a non-transparent way.
The primary goal of this paper is to deepen our understanding of sampling for $\ell^{2}$-regularized least squares approximations. Our main contribution is the identification of a norm inequality, analogous to that of unregularized least squares approximation (9), which fully characterizes the accuracy and conditioning of these methods. The following results are part of Theorem 7 and 8.
@@
A couple of important conclusions are in order.
  • The discretization condition (13) is a relaxation of (9); one can always find $A_{n,m}^{\epsilon}\geq A_{n,m}$. In order to see why, observe that both norm inequalities are equivalent for $\epsilon=0$. In other words: it is easier to satisfy (13) than (9) and, hence, regularization can only reduce the required amount of data. After formally introducing the concept of numerical redundancy, it will become clear that this relaxation is significant only for numerically redundant spanning sets. Note that we have reduced the two constants $\kappa^{\epsilon}_{M,N}$ and $\lambda^{\epsilon}_{M,N}$ introduced in [8, Theorem 1.3] to a single, more transparent constant.
  • The accuracy is now governed by (14) as opposed to (10). As discussed in section 2.2, this behaviour is expected when approximating with finite precision. As a result, the size of the expansion coefficients $\|\mathbf{x}\|_{2}$ is an important object of study, which was also one of the main conclusions of [6, 8].
  • Whereas (9) solely depends on the sampling operator $\mathcal{M}_{m}$ and the approximation space $V_{n}$, (13) also depends on the spanning set $\Phi_{n}$ and the regularization parameter $\epsilon$. These additional dependencies make its analysis significantly more complicated.
These results offer a new perspective on the influence of finite precision when working with redundant spanning sets. While finite precision leads to numerical regularization, which generally reduces accuracy compared to the best approximation, there is a benefit: regularization also decreases the amount of required information. Although predicting the exact interplay between these two effects remains challenging—particularly due to the dependence of (13) and (14) on the specific spanning set $\Phi_{n}$—this paper offers the tools to obtain a deeper understanding of practical approximation schemes.
@@ 1.4 Paper Overview
In section 2, we formalize the notion of numerical redundancy using ideas from low rank matrix theory and harmonic analysis. Our primary goal is to clearly identify when the effects discussed in this paper are significant. Using this framework we prove that the achievable error of a numerical approximation intrinsically depends on the achievable residual and the size of coefficients. In section 3, we motivate our focus on $\ell^{2}$-regularized approximation as a model for numerical approximation. Through a backward stability analysis we show that numerical algorithms fail to compute the expansion coefficients associated with the best approximation in the presence of numerical redundancy, yet succeed in computing $\ell^{2}$-regularized approximations. Furthermore, it is demonstrated that straightforward numerical orthogonalization cannot overcome the effects of finite-precision arithmetic. This implies that the results of this paper are relevant for a broad class of methods.
In the following part, we derive sampling theoretical results taking into account the effects of $\ell^{2}$-regularization. Section 4 formalizes how regularization relaxes the requirements for stable discretization, resulting in a decrease of the amount of data required for accurate function approximation. Our analysis is not limited to TSVD regularization, or even $\ell^{2}$-regularization; instead, we present a unified framework for regularized least squares problems with a general penalty term. In section 5, we show an additional advantage of the formulation (13): it enables constructive methods for selecting the sampling functionals $l_{j,m}$. Specifically, we identify an efficient random sampling distribution for $L^{2}$-approximations that takes into account the influence of regularization. Section 6 applies the results to two different settings: approximation that incorporates prior knowledge on the function to be approximated and approximation on irregular domains. In doing so, we obtain new results on sampling for Fourier extension frames. Specifically, we show that only a log-linear amount of uniformly random samples is sufficient for convergence down to machine precision in a Fourier extension frame. In contrast, an analysis that ignores the effects of regularization would suggest a need for quadratic oversampling.
2 Numerical redundancy
In this section, we aim to identify which spanning sets lead to significant rounding errors in the computation of expansion coefficients and how these errors affect the achievable accuracy. In order to do so, we introduce the concept of numerical redundancy. Notably, this analysis is independent of the specific discretization used.
2.1 Definition and links to frames theory
@@
i.e., $\mathbf{x}$ lies in the kernel of the synthesis operator $\mathcal{T}_{n}$ (5). When computing numerically, the notion of linear dependency changes drastically since computations can only be performed with a finite working precision $\operatorname{\epsilon_{\text{mach}}}$. A vector $\mathbf{x}$ is stored as $\bar{\mathbf{x}}$ where
@@
$\|\mathcal{T}_{n}\bar{\mathbf{x}}-\mathcal{T}_{n}\mathbf{x}\|_{H}=\|\mathcal{T}_{n}\Delta\mathbf{x}\|_{H}\leq\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}\|\mathbf{x}\|_{2}.$ (16)
@@
To take into account the influence of finite precision, it is common in low rank matrix theory to introduce a numerical rank. Whereas the rank of $\mathcal{T}_{n}$ equals $n$, i.e., the dimension of $V_{n}$, its $\epsilon$-rank can be defined as
@@
where the infimum is taken over all linear operators $\mathcal{X}_{n}$ mapping from $\mathbb{C}^{n^{\prime}}$ to $H$ [29]. The $\epsilon$-rank takes into account that, in the presence of rounding errors, the computed output $\mathcal{T}_{n}\bar{\mathbf{x}}$ behaves similarly as the exact output $\mathcal{X}_{n}\mathbf{x}$ of some linear operator $\mathcal{X}_{n}$ that lies close to $\mathcal{T}_{n}$. Indeed, similarly to (16) one has
$\|\mathcal{X}_{n}\mathbf{x}-\mathcal{T}_{n}\mathbf{x}\|_{H}\leq\epsilon\|\mathcal{T}_{n}\|_{2,H}\|\mathbf{x}\|_{2}$
@@
A spanning set $\Phi_{n}$ for the $n$-dimensional space $V_{n}$ is called “numerically redundant” if the $\epsilon$-rank of its synthesis operator $\mathcal{T}_{n}$ (17) is strictly less than $n$, where $\epsilon=\operatorname{\epsilon_{\text{mach}}}$ is the working precision associated with the numerical computations.
An elegant characterization of numerically redundant spanning sets follows from the field of harmonic analysis. To this end we introduce frames, which extend the concept of a Riesz basis to allow for the inclusion of linearly dependent sets. As defined by [19, Definition 1.1.1], the set $\Phi_{n}$ is a frame for $V_{n}$ if there exist frame bounds $0<A_{n}\leq B_{n}$ such that
@@
If we refer to the frame bounds $A_{n}$ and $B_{n}$ of $\Phi_{n}$ we will always assume that they are optimal, meaning that they are the largest and smallest constants that satisfy (18), respectively. $A_{n}$ and $B_{n}$ always exist, since every finite set is a frame for its span [19, Corollary 1.1.3]. Furthermore, they satisfy
@@
From the extremal properties of singular values [15, Lecture 23], it follows that
@@
where the minimum is taken over all linear operators $\mathcal{X}_{n}$ mapping from $\mathbb{C}^{n^{\prime}}$ to $H$. Combining this result with Definition 1, one can conclude that $\Phi_{n}$ is numerically redundant if and only if the smallest nonzero singular value of $\mathcal{T}_{n}$ is less than or equal to $\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}$. The final result follows from (19). ∎
@@
Based on the properties of $\Psi$, one can determine whether $\Phi_{n}$ is numerically redundant for sufficiently large $n$. In [6, section 4] this is analysed for a linearly independent infinite-dimensional frame $\Psi$. A set $\Psi$ is an infinite-dimensional frame for $H$ if there exist constants $A,B>0$ such that
@@
Furthermore, $\Psi$ is linearly independent if every finite subsequence is linearly independent [19, §3-4]. Note that in contrast to the finite-dimensional case, it does not hold that every infinite set is a frame for its closed span. Again, if we refer to the frame bounds $A$ and $B$ of $\Psi$ we assume that they are optimal. If $\Psi$ is a frame for $H$, analogously to (20) one has that for any $f\in H$ there exist coefficients $\mathbf{x}\in\ell^{2}(\mathbb{N})$ satisfying
@@
Assuming that $\Psi$ is a linearly independent frame, numerical redundancy can now be identified from analysing the upper and lower frame bound of $\Phi_{n}$ as $n$ increases. Following [6, Lemma 5], the upper frame bound $B_{n}$ behaves nicely; it is shown that $\{B_{n}\}_{n\in\mathbb{N}}$ is monotonically increasing and $B_{n}\to B$, as $n\to\infty$. However, the lower frame bound $A_{n}$ can behave wildly; it is shown that $\{A_{n}\}_{n\in\mathbb{N}}$ is monotonically decreasing and $\inf_{n}A_{n}=0$ if $\Psi$ is overcomplete, which means that there exist unit-norm coefficients $\mathbf{c}\in\ell^{2}(\mathbb{N})$ for which